# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 2.3.1. First variation and Flat structure
We now briefly recap the principal notions of differential calculus in Wasserstein spaces. For more details, see [AGS, Section  10.3] and [CarmonaDelarue, Chapter 5]. At the end of the discussion, we will select a unique differential that can be used for equations (1.1) and (1.2).
We first observe that the Wasserstein space can be seen as a convex subset of the vector space of finite Borel measures $\mathcal{M}(\mathbb{R}^{d})$. We prefer to give a more general definition inspired by the definitions of first derivative over $\mathcal{P}_{2}(\mathbb{R}^{d})$ contained in [PierreCaNotes] and [santambrogio2015optimal]. The growth assumption is taken by [CarmonaDelarue, Chapter  5].
@@
We say that $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is $C^{1}$ (resp. partially $C^{2}$) if $\mathcal{D}_{\mu}U:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ is jointly continuous with uniform quadratic growth (there exists $\frac{\partial}{\partial x}\mathcal{D}_{\mu}U$ that is jointly continuous and with uniform quadratic growth).
Since the convex interpolation is restricted to be in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\int\text{cost}(\nu-\mu)dy=0$, in supp($\mu$), the flat derivative $\mathcal{D}_{\mu}U$ is defined up to a constant in depending only on $\mu$. We select this constant s.t. $\int_{\mathbb{R}^{d}}\mathcal{D}_{\mu}U(\mu,x)d\mu=0.$ Thanks to this choice of the constant, we have the following representation formula: if $U$ is flat differentiable at $\mu$, then
(2.12) $\mathcal{D}_{\mu}U(\mu,y)=\lim_{h\to 0^{+}}\frac{1}{h}\left(U\left((1-h)\mu+h\delta_{y}\right)-U(\mu)\right),\quad\forall y\in\text{supp}(\mu).$
2.3.2. Wasserstein subdifferential calculus
Various notions of subdifferentiability have been introduced in the literature. Let us start with two classical notions that are given with two different points of view. The first one sees the elements of the differential as plans while the second one as vector valued functions. The two notions can be related via the barycentric projection map introducing an intermediate notion of differential.
In the following, we will provide only the definition of subdifferential of a proper and lower semicontinuous function since the superdifferential of a proper and upper semicontinuous function $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,\infty)$ can be obtained setting $\partial^{+}_{\cdot}\phi(\mu):=\big\{-1\cdot\gamma:\gamma\in\partial_{\cdot}^{-}(-\phi)(\mu)\big\}$ in the definitions 2.4 2.6, 2.7 and 2.9. As already remarked, this operation leaves invariant the geometric tangent space. Whenever $\partial^{-}_{\cdot}(-\phi)(\mu)\subset L^{2}_{\mu}(\mathbb{R}^{d})$, as in Definitions 2.5 and 2.8, we simply consider $\partial^{+}_{\cdot}\phi(\mu)=-\partial^{-}_{\cdot}(-\phi)(\mu)$. We denote by $\text{Dom}(\phi):=\Big\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}):\phi(\mu)<\infty\Big\}$
In what follows, we introduce the various notions of differentials within the formalism established above. For each definition, we explicitly indicate the corresponding reference.
Definition 2.4 (See [GigliPhd], Definition 5.14).
@@ 2.3.2. Wasserstein subdifferential calculus
Note that in [GigliPhd, Definition 5.14], the author additionally requires that $\partial^{-}_{K}\phi(\mu)\subset\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In the present work, we choose not to impose this restriction at the level of the definition, in order to allow for a more refined comparison with other notions.
Definition 2.5.
See [GANGBO2019119], Definition 3.1 ] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $p\in L^{2}_{\mu}(\mathbb{R}^{d})$ belongs to the Monge subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{M}\phi(\mu)$, if for all $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
@@
In order to formalize the link between the previous two classical notions, let us introduce the following intermediate definition, given in the spirit of Defintion 2.9 below, introduced by Bertucci in [BSOT].
@@
Note that the set $\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the subdifferential set defined in [DS].
The unique element in $\partial^{+}_{K}\phi(\mu)\cap\partial^{-}_{K}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, if any, is called the Wasserstein differential of $\phi$, see [GigliPhd, Definition  5.14]. In this case, we say that $\phi$ is differentiable at $\mu$ and $\partial_{\mu}\phi$ is the Wasserstein differential at $\mu$.
A stronger notion of subdifferentials can be obtained by requiring, in Definitions 2.4,2.5 and 2.6, that the corresponding inequalities hold for a larger class of perturbations, e.g. without assuming the measure vector field $\xi$ to lie in the tangent space. This restriction will select more stable elements in the subdifferentials, and it has already been used in the gradient flow theory developed in[AGS] to treat the Euler equation of the minima of the Yosida approximation of a given functional. We also remark that this strongest notion is linked with the so-called Hilbertian lift. See [BSOT, Remark  2.5 &\& Proposition  2.6] and the introduction in [GANGBO2019119].
Definition 2.7.
[See [AGS, Definition  10.3.1] Strong Fréchet Subdifferential] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu}(\gamma,\xi)$,
@@
Definition 2.9.
[See [BSOT], Definition 2.2] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the map strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S,I}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
@@
The above definition was introduced in [BSOT] to prove the well-posedness of viscosity solutions of Hamilton-Jacobi equation in $\mathcal{P}(\mathbb{T}^{d})$.
@@
Let us note that w.r.t. [BSOT, Definition  2.2], we preferred here to describe the elements of the subdifferential as plans instead of using the equivalent identification of $\gamma$ with $\gamma_{x}$ $\mu$-a.e. determined by $\gamma(dx,dy)=\gamma_{x}(dy)\mu(dx)$. We will do the same for the notion of viscosity solutions adopted in the first-order case.
@@
(Correspondent of [GANGBO2019119, Theorem  3.14] in our setting) Let $\phi:\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be a proper lower semicontinuous functional, and $\mu\in\text{Dom}(\phi)$. Then,
@@
In view of the inclusion $\partial^{-}_{S,M}\phi(\mu)\subseteq\partial^{-}_{M}\phi(\mu)$, it is sufficient to prove only the $\supseteq$ part. The proof is an adaptation of [Erbar2010, Lemma  3.2].
@@
By combining the two inequality we deduce that
@@
are both bounded by construction. We can then extract a, not relabeled, weak limit in $L^{2}_{\mu}(\mathbb{R}^{d})$ of the two sequences $\left(\frac{\text{Pr}(\widetilde{\xi}_{n})}{\varepsilon_{n}}\right)_{n},\left(\frac{\text{Pr}(\xi_{n})}{\varepsilon_{n}}\right)_{n}$, call them $\widetilde{\xi},{\xi}\in L^{2}_{\mu}(\mathbb{R}^{d})$, respectively. In particular,
@@
Consider now the $p\in\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ given at the beginning of the proof: since $W_{\mu}((\mathrm{Id}\times p)_{\mu},(\mathrm{Id}\times\nabla\psi)_{\mu})=\|p-\nabla\psi\|_{L^{2}_{\mu}(\mathbb{R}^{d})}$, a density argument and the alternative description of the tangent space given in (2.11) lead us to
@@ 2.4. Selection of the differential
Note that the intersection $\partial^{-}\phi(\mu)\cap\partial^{+}\phi(\mu)$ contains at most one element since $\partial^{+}_{S}\phi(\mu)\cap\partial^{-}_{S}\phi(\mu)$ does. Whenever this element exist, we say that the continuous function $\phi$ is differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. This element, denoted with $\partial_{\mu}\phi$, coincides with the Wasserstein gradient (and it is always a map, otherwise $\text{Pr}(\partial_{\mu}\phi)$ would be a different element in the intersection).
We end by stressing that thanks to Lemma (2.12), at absolutely continuous measures, our notion of subdifferential coincides with the one adopted in [DS], up to identification via the barycentric projection, in the definition of viscosity solutions of (1.1). Moreover, being $\partial^{-}\subseteq\partial^{-}_{S,I}$, our subdifferential is a subset of the subdifferential introduced in [BSOT] to treat viscosity solutions of (1.2).
@@
A similar result which explores the link between extrinsic and intrinsic geometry has been obtained in the recent paper [CavSavSod23], where an equivalence between geodesic convexity (intrinsic convexity) and strong (or total) convexity (extrinsic convexity) has been established for continuous functions, see Theorem 9.1 therein.
We state here an important result (See [BSOT, Proposition 2.4]) that it is fundamental in comparison arguments and describes the superdifferentiability of the Wasserstein distance squared that we need. [^4] Actually, [BSOT, , Proposition 2.4] is stated for $\partial^{+}_{S,I}$. However, a careful inspection of the simple proof reveals that the very same argument gives the result stated here for our choice of superdifferential.
Proposition 2.14.
@@ 2.5. Main Hypotheses
    $|H(x,p,\mu)-H(y,q,\nu)|\leq C\bigl(1+|p|+|q|\bigr)\bigl(W_{2}(\mu,\nu)+|x-y|+|p-q|\bigr).$
@@
The diffusion term in (1.2) is of the form $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$

### v2

@@ 2.3.1. First variation and Flat structure
We now briefly recap the principal notions of differential calculus in Wasserstein spaces. For more details, see [2, Section  10.3] and [14, Chapter 5]. At the end of the discussion, we will select a unique differential that can be used for equations (1.1) and (1.2).
We begin with the definition of the flat derivative, noting that the Wasserstein space can be seen as a convex subset of the vector space of finite Borel measures $\mathcal{M}(\mathbb{R}^{d})$. We adopt a more general definition, inspired by the notions of first derivative on $\mathcal{P}_{2}(\mathbb{R}^{d})$ introduced in [12] and [45]. The growth assumption is taken from [14, Chapter  5].
@@
We say that $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is $C^{1}$ (resp. fully $C^{1}$) if there exists $\mathcal{D}_{\mu}U:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}$ that is jointly continuous with uniform quadratic growth (resp. if $U$ is $C^{1}$ and there exists $\nabla\mathcal{D}_{\mu}U:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that is jointly continuous and with uniform linear growth).
We say that $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is partially $C^{2}$ if $U$ is fully $C^{1}$ and there exists $\nabla^{2}\mathcal{D}_{\mu}U:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to M_{d\times d}(\mathbb{R})$ that is jointly continuous and uniformly bounded.
Since the convex interpolation is restricted to be in $\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\int\text{cost}(\nu-\mu)dy=0$, the flat derivative $\mathcal{D}_{\mu}U$ is defined up to a constant depending only on $\mu$. We select this constant s.t. $\int_{\mathbb{R}^{d}}\mathcal{D}_{\mu}U(\mu,x)d\mu(x)=0.$ Due to this choice of constant, we have the following representation formula: if $U$ is flat differentiable at $\mu$, then
(2.14) $\mathcal{D}_{\mu}U(\mu,y)=\lim_{h\to 0^{+}}\frac{1}{h}\left(U\left((1-h)\mu+h\delta_{y}\right)-U(\mu)\right),\quad\forall y\in\mathbb{R}^{d}.$
2.3.2. Wasserstein differential calculus
Various notions of differentiability have been introduced in the literature. Let us start with two classical notions that are given with two different points of view. The first one sees the elements of the differential as plans while the second one as vector valued functions. The two notions can be related via the barycentric projection map introducing an intermediate notion of differential.
In the following, we will provide only the definition of subdifferential of a proper and lower semicontinuous function since the superdifferential of a proper and upper semicontinuous function $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to[-\infty,\infty)$ can be obtained by setting $\partial^{+}_{\cdot}\phi(\mu):=\big\{-1\cdot\gamma:\gamma\in\partial_{\cdot}^{-}(-\phi)(\mu)\big\}$ in the definitions 2.4 2.6, 2.7 and 2.9. As already remarked, this operation leaves invariant the geometric tangent space. Whenever $\partial^{-}_{\cdot}(-\phi)(\mu)\subset L^{2}_{\mu}(\mathbb{R}^{d})$, as in Definitions 2.5 and 2.8, we simply consider $\partial^{+}_{\cdot}\phi(\mu)=-\partial^{-}_{\cdot}(-\phi)(\mu)$. We denote by $\text{Dom}(\phi):=\Big\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}):\phi(\mu)<\infty\Big\}$ the effective domain of $\phi$.
In what follows, we introduce the various notions of differential within the formalism established above. For each definition, we explicitly indicate the corresponding reference.
Definition 2.4 (See [37], Definition 5.14).
@@ 2.3.2. Wasserstein differential calculus
Note that in [37, Definition 5.14], the author additionally requires that $\partial^{-}_{K}\phi(\mu)\subset\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$. In the present work, we choose not to impose this restriction at the level of the definition, in order to allow for a more refined comparison with other notions.
Definition 2.5 (See [36], Definition 3.1).
Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $p\in L^{2}_{\mu}(\mathbb{R}^{d})$ belongs to the Monge subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{M}\phi(\mu)$, if for all $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$
@@
In order to formalize the link between the previous two classical notions, let us introduce the following intermediate definition, given in the spirit of Definition 2.9 below, introduced in [9].
@@
Note that the set $\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the subdifferential set defined in [30].
The unique element in $\partial^{+}_{K}\phi(\mu)\cap\partial^{-}_{K}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, if any, is called the Wasserstein differential of $\phi$, see [37, Definition  5.14]. In this case, we say that $\phi$ is differentiable at $\mu$ and $\partial_{\mu}\phi$ is the Wasserstein differential at $\mu$.
A stronger notion of subdifferential can be obtained by requiring, in Definitions 2.4,2.5 and 2.6, that the corresponding inequalities hold for a larger class of perturbations, e.g. without assuming the measure vector field $\xi$ to lie in the tangent space. This restriction will select more stable elements in the subdifferentials, and it has already been used in the gradient flow theory developed in[2] to treat the Euler equation of the minima of the Yosida approximation of a given functional. We also remark that this strongest notion is linked with the so-called Hilbertian lift. See [9, Remark  2.5 &\& Proposition  2.6] and the introduction in [36].
Definition 2.7 (See [2], Definition 10.3.1).
Strong Fréchet Subdifferential] Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ and $\theta\in\Gamma_{\mu}(\gamma,\xi)$,
@@
Definition 2.9 (See [9], Definition 2.2).
Let $\phi\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be proper and lower semicontinuous. Let $\mu\in\text{Dom}(\phi)$. We say that $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ belongs to the map strong subdifferential of $\phi$ at $\mu$, denoted $\partial^{-}_{S,I}\phi(\mu)$, if for all $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
@@
The above definition was introduced in [9] to prove the well-posedness of viscosity solutions of Hamilton-Jacobi equations in $\mathcal{P}(\mathbb{T}^{d})$.
@@
Let us note that w.r.t. [9, Definition  2.2], we preferred here to describe the elements of the subdifferential as plans instead of using the equivalent identification of $\gamma$ with $\gamma_{x}$ $\mu$-a.e. determined by $\gamma(dx,dy)=\gamma_{x}(dy)\mu(dx)$. We do the same for the notion of viscosity solutions adopted in the first-order case.
@@
(Correspondent of [36, Theorem  3.14] in our setting) Let $\phi:\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ be a proper lower semicontinuous functional, and $\mu\in\text{Dom}(\phi)$. Then,
@@
In view of the inclusion $\partial^{-}_{S,M}\phi(\mu)\subseteq\partial^{-}_{M}\phi(\mu)$, it is sufficient to prove only the $\supseteq$ part. The proof is an adaptation of [31, Lemma  3.2].
@@
By combining the two inequalities we deduce that
@@
are both bounded by construction. We can then extract a weak limit in $L^{2}_{\mu}(\mathbb{R}^{d})$ of the two, not relabeled, sequences $\left(\frac{\text{Pr}(\widetilde{\xi}_{n})}{\varepsilon_{n}}\right)_{n},\left(\frac{\text{Pr}(\xi_{n})}{\varepsilon_{n}}\right)_{n}$, that we call $\widetilde{\xi},{\xi}\in L^{2}_{\mu}(\mathbb{R}^{d})$, respectively. In particular,
@@
Consider now the $p\in\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ given at the beginning of the proof: since $W_{\mu}((\mathrm{Id}\times p)_{\#}{\mu},(\mathrm{Id}\times\nabla\psi)_{\#}{\mu})=\|p-\nabla\psi\|_{L^{2}_{\mu}(\mathbb{R}^{d})}$, a density argument and the alternative description of the tangent space given in (2.13) lead us to
@@
Remark 2.13.
In view of the previous lemma, we will frequently exploit the identification provided by the isometric relation (2.15). In particular, we shall not distinguish between the measure $(\mathrm{Id}\times p)_{\#}\mu\in\partial^{-}_{S}\phi(\mu)\cap\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ and its projection $p:=\mathrm{Pr}\big((\mathrm{Id}\times p)_{\#}\mu\big),$ which is regarded as an element of $\partial^{-}_{M}\phi(\mu)\cap\mathrm{Tan}^{M}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@ 2.4. Selection of the differential
Note that the intersection $\partial^{-}\phi(\mu)\cap\partial^{+}\phi(\mu)$ contains at most one element since $\partial^{+}_{S}\phi(\mu)\cap\partial^{-}_{S}\phi(\mu)$ does. Whenever this element exists, we say that the continuous function $\phi$ is differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. This element, denoted with $\partial_{\mu}\phi$, coincides with the Wasserstein gradient (and it is always a map, otherwise $\text{Pr}(\partial_{\mu}\phi)$ would be a different element in the intersection). As a consequence, if $\phi$ is fully $C^{1}$ then $\nabla\mathcal{D}_{\mu}\phi=\partial_{\mu}\phi$, see the structural results [14, Lemma  5.61 & Theorem  5.65].
We end by stressing that thanks to Lemma 2.12, at absolutely continuous measures, our notion of subdifferential coincides, up to identification via the barycentric projection, with the one adopted in [30] in the definition of viscosity solutions of (1.2). Moreover, being $\partial^{-}\subseteq\partial^{-}_{S,I}$, our subdifferential is a subset of the subdifferential introduced in [9] to treat viscosity solutions of (1.1).
@@
A similar result which explores the link between extrinsic and intrinsic geometry has been obtained in the recent paper [15], where an equivalence between geodesic convexity (intrinsic convexity) and strong (or total) convexity (extrinsic convexity) has been established for continuous functions, see Theorem 9.1 therein.
We state here an important result that is fundamental in comparison arguments and describes the superdifferentiability of the Wasserstein distance squared that we need. [^3] Actually, [9, Proposition 2.4] is stated for $\partial^{+}_{S,I}$. However, a careful inspection of the simple proof reveals that the very same argument gives the result stated here also for our choice of superdifferential.
Proposition 2.15 (See [9] Proposition 2.4).
@@
We remark that a complete characterization of the elements in the superdifferential of the Wasserstein distance squared is given in [4, Theorem.  5.1.19].
@@ 2.5. Main Hypotheses
    $|H(x,p,\mu)-H(y,q,\nu)|\leq C\bigl(1+|p|+|q|\bigr)\bigl(W_{2}(\mu,\nu)+|x-y|+|p-q|\bigr);$
@@
The diffusion term in (1.2) is of the form $a(\mu,x)=\varepsilon\text{Id}_{d\times d}$.
2.6. Weak notion of convergence
The space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is not locally compact, since the underlying space $\mathbb{R}^{d}$ is non-compact (see [2, Rmk. 7.1.9]). This creates difficulties in the application of the doubling variable technique, where one needs to extract a maximizing sequence; see Lemma 4.3. For this reason, we introduce suitable continuity assumptions on the functions under consideration.
Definition 2.17.
We say that a sequence $(\mu_{n})_{n}\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ weakly converges in $\mathcal{P}_{2}(\mathbb{R}^{d})$ to $\mu\in\mathcal{P}(\mathbb{R}^{d})$ if $\mu_{n}$ converges narrowly to $\mu$ and the second moments are uniformly bounded, namely
$\sup_{n}\mathcal{M}_{2}(\mu_{n})<\infty.$
Note that, by the lower semicontinuity of the second moment with respect to narrow convergence, we have
(2.18) $\mathcal{M}_{2}(\mu)\leq\liminf_{n\to\infty}\mathcal{M}_{2}(\mu_{n})<\infty,$
for every sequence $(\mu_{n})_{n}$ converging to $\mu$ in this topology. In particular, this ensures that $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. In general, the inequality in (2.18) is strict; when equality holds, one has $W_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$, see [50, Theorem 6.9].
By the Prokorov Theorem [2, Theorem  5.1.3], the space $\mathcal{P}_{2}(\mathbb{R}^{d})$, endowed with this topology, is locally compact. [^4] This follows from the tightness estimate $\mu_{n}\bigl(\mathbb{R}^{d}\setminus B_{R}(0)\bigr)\leq\frac{1}{R^{2}}\,\mathcal{M}_{2}(\mu_{n}),\qquad\forall\,R>0.$
Definition 2.18.
We say that a function $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}\cup\{+\infty\}$ is weakly lower semicontinuous in $\mathcal{P}_{2}(\mathbb{R}^{d})$ if it is lower semicontinuous with respect to the topology defined by the weak convergence in Definition 2.17.
We also observe that any function which is lower semicontinuous in $\mathcal{P}_{p}(\mathbb{R}^{d})$ for some $1\leq p<2$ is weakly lower semicontinuous in $\mathcal{P}_{2}(\mathbb{R}^{d})$.
Examples of functions that are lower semicontinuous with respect to this notion of convergence include the Wasserstein distance (see [2, Prop. 7.1.3]) and the entropy functional (see [50, Theorem 5.26], where lower semicontinuity follows directly from its variational representation).
We conclude this subsection by emphasizing that, when one works on $\mathcal{P}_{2}(\mathbb{T}^{d})$, weak convergence and $W_{2}$ convergence coincide.